### v1

# Slamming: Training a Speech Language Model on One GPU in a Day

Abstract
We introduce Slam, a recipe for training high-quality Speech Language Models (SLMs) on a single academic GPU in 24 hours. We do so through empirical analysis of model initialisation and architecture, synthetic training data, preference optimisation with synthetic data and tweaking all other components. We empirically demonstrate that this training recipe also scales well with more compute getting results on par with leading SLMs in a fraction of the compute cost. We hope these insights will make SLM training and research more accessible. In the context of SLM scaling laws, our results far outperform predicted compute optimal performance, giving an optimistic view to SLM feasibility. See code, data, models, samples - https://pages.cs.huji.ac.il/adiyoss-lab/slamming.
Slamming: Training a Speech Language Model on One GPU in a Day
Gallil Maimon*, Avishai Elmakies*, Yossi Adi Equal Contribution The Hebrew University of Jerusalem gallil.maimon@mail.huji.ac.il
1 Introduction
Speech Language Models (SLMs) have gained significant interest from researchers Peng et al. (2024a); Cui et al. (2024); Ji et al. (2024); Latif et al. (2023), demonstrating remarkable performance in traditional speech tasks Wang et al. (2023); Elmakies et al. (2025), diverse generative applications Yang et al. (2023, 2024b), and reasoning over speech and audio signals Tang et al. (2024); Chu et al. (2023).
SLMs can generally be classified into two main categories: (i) generative speech Language Models (LMs) (which can also incorporate text) and (ii) speech-aware LMs. The first category follows a similar pre-training approach to text-based Large Language Models (LLMs), directly maximising the likelihood of speech considering both input and output, typically by representing audio as a sequence of discrete tokens. The second category consists of pre-trained text LMs adapted to process speech inputs. In this work, we focus on the first.
Figure 1: Comparing Topic-StoryCloze performance of different SLMs as a function of training compute. Model size is indicated by the size of the circle.
Training high-quality SLMs can be highly resource intensive Hassid et al. (2024); Cuervo and Marxer (2024); Zeng et al. (2024); Nguyen et al. (2025); Défossez et al. (2024). For example,  Nguyen et al. (2025) trained their SLM on approximately $570k$ hours of speech data, while  Défossez et al. (2024) utilised around $7M$ hours. Additionally,  Cuervo and Marxer (2024) proposed SLM scaling laws, suggesting that training high-quality SLMs requires $\sim 3X$ more data compared to text-based counterparts. These computational demands restrict the required fundamental research aimed at enhancing SLMs, such as advancements in speech tokenisation, efficient acoustic modelling, etc.
In the Natural Language Processing (NLP) community, numerous studies have investigated efficient model training techniques, including masked language models such as Cramming (Geiping and Goldstein, 2023) and ModernBERT (Warner et al., 2024), along with next-token prediction LLMs such as MobileLLM (Liu et al., 2024b). These methods include implementation efficiencies, architectural improvements, data selection strategies, and enhancements to the overall training pipeline.
Inspired by Cramming Geiping and Goldstein (2023) in text, we investigate compute-limited SLM training, which we term Slamming. We pose the question: Is it possible to train high-quality SLMs using a single GPU within 24 hours? For that, we conduct an extensive empirical analysis exploring how different training components influence performance. From this, we derive a training recipe that maximises model performance within a fixed compute budget. Specifically, we investigate the impact of model initialisation and architecture, various optimisers and learning rate schedulers, data selection strategies - including the role of synthetic data, text-interleaving and preference optimisation.
We believe that developing these training strategies and proving their feasibility will empower the speech and audio research community to advance SLMs beyond the scope of large, well-funded academic and industrial labs. Figure 1 illustrates the performance of various SLMs relative to their training compute budget, with circle sizes representing the size of the models. Furthermore, we compare our results with the scaling performance predicted from Cuervo and Marxer (2024). Although the authors present a somewhat pessimistic view of the computational resources needed to train high-quality SLMs, we empirically show that reality is more promising, demonstrating that it is possible to significantly exceed the predicted performance per unit of compute. We encourage the community to refine and expand scaling laws specifically tailored for SLM training across various settings.
Our main contributions are:
  1. 1.
    We introduce Slam, a training recipe for efficiently training high-quality SLMs using a single $A5000$ GPU within $24$ hours.
  2. 2.
    We carry out extensive experiments exploring model initialisation and architecture, optimisation, data collection and generation, and training objectives (i.e., preference optimisation and text-speech interleaving), providing insights into the impact of each component on model performance.
  3. 3.
    Building on these insights, we scale the compute budget to two $A100$ GPUs for $48$ hours and demonstrate that our model achieves performance on par with state-of-the-art models that require substantially more compute.
We open-source all code, models, training recipes, and synthetic datasets.
2 Related Work
Efficient training. Enhancing the efficiency of neural network training has been extensively studied (Shen et al., 2023). Hajimolahoseini et al. (2023); Wang et al. (2024) examined the impact of data selection on Large Language Model (LLM) training and introduced efficient data selection methods. Muhamed et al. (2024) proposed using structured sparse gradients to enhance compute efficiency in LLM training, while Rawat et al. (2024) explored the potential of leveraging smaller language models to improve the training efficiency of larger LLMs. Lv et al. (2024) investigated the use of low-dimensional projections for attention parameters to enhance training efficiency. Meanwhile, Neiterman and Ben-Artzi (2024) proposed applying LayerDrop as a technique to optimise neural network training.
More closely related to our work, Li et al. (2023) propose a training strategy for developing LLMs within a $100k\$$ budget. Warner et al. (2024) introduce ModernBERT, an efficient training pipeline for optimising BERT models, while Izsak et al. (2021) outline a method for training a BERT model in $24$ hours using $8$ GPUs. The most relevant work to ours is Cramming Geiping and Goldstein (2023), where the authors conduct an in-depth analysis of masked LM training on a single GPU in one day.
While these studies offer valuable insights, they primarily focus on training text models, such as LLMs and masked LMs. In the speech domain, similar research has been conducted on self-supervised representation models (Liu et al., 2024a), but not on SLMs. In this work, we address this gap by focusing on efficient SLM training.
Generative speech language models were explored under various setups (Lakhotia et al., 2021; Kharitonov et al., 2021). Lakhotia et al. (2021) were the first to show how raw and uncurated speech data can be leveraged into building a Generative Spoken Language Modeling (GSLM) system. Next, Borsos et al. (2023) proposed a cascade version using both coarse and fine speech tokens. Such a modelling framework opened up a new and promising research direction for processing and modelling spoken data, such as speech resynthesis (Polyak et al., 2021), speaking style conversion (Kreuk et al., 2021; Maimon and Adi, 2023), dialogue modelling Nguyen et al. (2022), speech-to-speech translation (Popuri et al., 2022; Peng et al., 2024b), etc. Nachmani et al. (2024) proposed augmenting a text Language Model (LM) with continuous speech data to improve spoken question-answering tasks. Recently, Park et al. (2024) proposed SLM based on state-space models (Gu et al., 2021) to further push long context-efficient modelling, while  Lin et al. (2024) proposed to fine-tune SLMs using direct preference optimisation (Rafailov et al., 2024) obtained from text LLM rankings.
Similar to text LLMs, training SLMs often demands large-scale datasets. For instance, Moshi Défossez et al. (2024) was trained on $7$ million hours of speech data, SpiritLM Nguyen et al. (2025) utilized $560k$ hours, and TWIST Hassid et al. (2024) was trained on approximately $150k$. Recently, Cuervo and Marxer (2024) introduced the first scaling laws for SLMs, suggesting that achieving comparable performance to text LMs requires three times more tokens. In this work, we focus on reducing the computational demands while maintaining performance comparable to leading SLMs.
3 Setup
In this study, we explore decoder-only generative SLMs, which aim at maximising the likelihood of speech samples represented as discrete tokens. We examine both purely speech-based SLMs trained on speech tokens and joint speech-text SLMs using interleaving strategies (Nguyen et al., 2025). Similarly to  Hassid et al. (2024); Lakhotia et al. (2021), we obtain speech tokens by quantising continuous latent representations of a self-supervised speech representation model using the k-means algorithm, often known as semantic tokens. Specifically, we utilise a multilingual HuBERT (Hsu et al., 2021) model running at $25$ Hz, as employed in  Hassid et al. (2024). We then train SLMs by minimising the negative log-likelihood of the input segments.
Unless mentioned otherwise, all SLMs are trained using a single $A5000$ GPU ($24GB$ VRAM) along with $16$ CPU cores for $24$ hours. We deliberately focus on this constrained compute budget, assuming that most academic labs can access similar resources, thereby ensuring the accessibility of our research. The training data is pre-processed, i.e. extracting HuBERT units and dividing data into chunks, and stored prior to model training. As a result, this pre-processing time is excluded from the compute budget. This approach, aligned with Geiping and Goldstein (2023), is practical since many research experiments utilise the same pre-processed data. We additionally do not count the time for running validation and visualisations as they are not used as part of the optimisation pipeline and only used for demonstration purposes.
Evaluation metrics. We assess all SLMs using four distinct evaluation metrics. The first three are based on likelihood evaluation, while the fourth is a generative metric. For likelihood based modelling we consider sBLIMP Dunbar et al. (2021), Spoken Story Cloze (sSC)), and Topic Story-Cloze (tSCHassid et al. (2024). For modelling-likelihood metrics, we evaluate the likelihood assigned by the SLMs to pairs of speech utterances, consisting of a positive example and a distractor. We calculate the percent of pairs in which the SLM assigns higher likelihood to the positive sample. sBLIMP focuses on grammatical abilities thus the negative is ungrammatical version of the positive. sSC and tSC focus on semantic modelling abilities. In sSC, the distractor suffix is taken from the original textual StoryCloze dataset (Mostafazadeh et al., 2016), allowing to assess fine-grained semantic speech understanding. In tSC, however, the distractor suffix is drawn from a different topic, enabling us to evaluate the model’s ability to understand the overall semantic concept.
Finally, to assess the generative abilities of SLMs, we compute generative perplexity (GenPPL). Following the approach of  Lakhotia et al. (2021); Hassid et al. (2024), we provide the SLM with a short speech prompt and generate speech tokens continuation. We use unit-vocoder with duration prediction to convert the tokens into speech (Polyak et al., 2021; Hassid et al., 2024). The generated speech is then transcribed, and its Perplexity (PPL) is evaluated using a pre-trained text LLM. To minimise the impact of token repetition on PPL measurements, we ground the generated text using diversity metrics derived from the auto-BLEU score (Lakhotia et al., 2021). Similarly to  Lin et al. (2024) we use bigram auto-BLEU. In other words, we ensure that all models achieve similar auto-BLEU scores, allowing for a fair comparison of PPL. Specifically, we transcribe speech segments using Whisper-large-v3333-turbo model (Radford et al., 2023) and measure PPL using Llama-$3.2$-$1$B model (LLama, 2024).
Software efficiency. To maximise performance within $24$ hours of model training, we leverage multiple efficient implementations. Through extensive performance testing, we found that using bfloat16161616  Kalamkar et al. (2019) alongside FlashAttention2222 (Dao, 2023) and data packing provided the most efficient compute performance in our setup. We also experimented with model compilation using torch.compile (Ansel et al., 2024), but it lacked native compatibility with FlashAttention2222 at the time of our study, and its performance without FlashAttention2222 was subpar. Future work could investigate this further with more efficient attention implementations Shah et al. (2024); Li et al. (2024).
To enable rapid and scalable experimentation, we developed a specialised library for SLM training that supports various model architectures, training objectives, and evaluation metrics. This library accommodates both TWIST-style training, text-speech interleaving, preference optimisation, etc. We will open-source this package along all models weights and training recipes, aiming to empower the community to further explore SLMs.
4 Investigations
With this setup, we systematically analyse and ablate each component of the training pipeline, ultimately refining an optimised cook-book for training SLMs. We specifically examine the influence of model family, initialisation, size, and architectural choices (e.g., dropout, positional embedding, etc.). We analyse optimisation parameters and data characteristics. Lastly, we explore alternative training objectives beyond standard next-token prediction, including speech-text interleaving and direct preference optimisation using synthetic data.
4.1 Model & Optimisation
Hyper-parameters. Unless specified otherwise, we use a context length of $512$ tokens and an effective batch size of $256$, employing gradient accumulation when necessary, as preliminary results indicated this configuration yields the best overall performance. We set the peak learning rate to $31e-$ 3 to enhance training speed and use a warmup period of $1\%$ of the total training steps, as this proved more effective than the fixed $100$-step warmup used in the original TWIST. To improve training stability, particularly with large learning rates, we apply gradient normalisation with a norm of $0.5$ at no additional cost, following Geiping and Goldstein (2023). Unless modified later in our investigation, we use an inverse-square root scheduler and the AdamW optimiser (Loshchilov, 2017).
Figure 2: Comparing PPL of different models of similar parameter count, with and without TWIST initialisation.
Initialisation. Hassid et al. (2024) empirically demonstrated that initialising SLMs with pre-trained text LMs can enhance convergence speed and improve model performance. We examine the effect of this initialisation within our setup across different model types. To do so, we train multiple models, both with and without TWIST initialisation, while staying within our compute budget. As shown in Figure 2, TWIST initialisation benefits all evaluated models at the beginning of training, though its overall impact by the end varies. Notice, the x-axis in Figure 2 represents theoretical FLOPs, calculated as $6*N_{params}*D_{tokens}$ following Hoffmann et al. (2022). However, due to variations in model architecture and implementation, practical efficiency differs, leading to varying amounts of compute processed within $24$ hours.
Results suggest that benefits of TWIST initialisation can be substantial, especially for top-performing models like Qwen2.52.52.52.5. As a result, we prioritise investigations based on existing pre-trained text LMs. Interestingly, the results in Figure 2 demonstrate that Qwen2.52.52.52.5 outperforms other models even without TWIST initialisation, perhaps suggesting that their architectural design choices or size might also provide some benefit.
Optimal model size & family. Cuervo and Marxer (2024) conducted a scaling analysis on GSLM-style SLMs, estimating the optimal model size and token count for a compute-efficient model. However, using a text LM initialisation might impact these findings. As we observe, TWIST initialisation greatly impact model performance, suggesting that prioritising larger models may be more effective than simply increasing the dataset size. Additionally, various model families gain different advantages from TWIST initialisation; for example, Qwen2.52.52.52.5 models show significantly better performance compared to OPT models. In Figure 3, we compare the results under the pre-defined compute budget within model families [^1] We use the text LM original names for clarity, but note that the actual size will be notably smaller due to reduced vocabulary size, e.g Qwen2.52.52.52.5-$0.5$B has $358$M parameters. Full model sizes can be found in Appendix B.. We note that the best model sizes for both MobileLLM Liu et al. (2024b), SmolLM2222 Allal et al. (2025) and Pythia Biderman et al. (2023) are $\sim 300M$ parameters, while for OPT the best is $125$M. According to Cuervo and Marxer (2024), the estimated optimal model size is approximately $66$M parameters. However, the best-performing model, Qwen2.52.52.52.5, is significantly larger. Since there are no smaller models in this family, it is difficult to determine whether this deviation is due to the quality of the initialisation or other factors. Moving forward, we proceed with both OPT-$125$M and Qwen2.52.52.52.5-$0.5$B.
Figure 3: Comparing PPL of different models under TWIST initialisation.
Dropout. The original OPT models includes dropout to mitigate overfitting. Although dropout is beneficial for regularisation, it effectively decreases the number of gradient updates per parameter without shortening the update-step wall time. Hence, reduces the number of parameter updates per second. Following Geiping and Goldstein (2023), we experiment with removing dropout and observed improved performance in our setup.
Positional Encoding. Transformers rely on positional encoding to capture the order of input tokens. Many modern LMs, including the Qwen models, use Rotary Position Embedding (Su et al., 2024). This method uses a hyperparameter, $\theta$, to control the trade-off between granularity and the ability to handle long contexts. $\theta$ is often tuned to accommodate longer context lengths (Yang et al., 2024a; Roziere et al., 2023). Since our context length is significantly shorter than that of the original LLM, we explore reducing $\theta$ for potential performance gains. Our findings show that setting $\theta=10,000$ with a context length of $1024$ enhances performance, so we adopt this configuration moving forward. We note that since we increase the context length, we need to reduce the batch size as well, to not run into memory problems when training. We reduce the batch size by a half and keep the same amount of gradient accumulation steps, which gives us an effective batch size of $128$.
Figure 4: Comparing validation PPL of our best model with different optimisers and schedulers.
Optimiser and Scheduler. Various optimisers and schedulers have been developed to enhance training efficiency, reduce memory usage (Shazeer and Stern, 2018; Dettmers et al., 2022), or accelerate convergence (Pagliardini et al., 2024; Chen et al., 2023). With limited compute, faster convergence and better memory efficiency can be especially important. We first consider efficient optimisers, specifically AdamW with fused kernels, and $8$-bit AdamW, but observe no notable improvements in batch size or runtime compared to standard AdamW. This could do with the relatively small model size, resulting in a minimal memory footprint of the optimisers. We then compare AdamW with two state-of-the-art optimisers: AdaLomo Lv et al. (2023) and AdEMAMeix (Pagliardini et al., 2024). Results, presented in Figure 4, suggest that with the original InverseSqrt scheduler used by Hassid et al. (2024), using AdEMAMeix improves validation loss, compared to AdamW, with AdaLomo far behind.
Next, we analyse a cosine decay learning rate scheduler, in place of the original InverseSqrt as this was shown to improve convergence Loshchilov and Hutter (2016). We consider the previous optimisers, and provide the validation loss throughout training in Figure 4. We see that this notably improved the loss for AdamW, and slightly harmed results for AdEMAMeix. Overall, AdamW with a cosine schedule provide the best setup, far outperforming the original setup.
4.2 Data
Next, we examine how the training data-mix influences performance in a compute-constrained setting. Specifically, we explore whether diversity in accents, speaking styles, etc. is beneficial and assess whether synthetic data can enhance semantic modelling abilities.
Model Data Metric
[…]
OPT125M 55.28 55.46 75.18 96.8
55.06 55.00 74.83 116.6
55.88 54.52 70.82 160.3
55.65 54.78 70.18 172.7
Qwen-0.5B 56.45 55.59 78.01 88.3
56.17 55.37 77.13 101.3
56.60 53.50 71.14 145.4
56.10 53.72 70.66 161.8
Table 1: Analysing impact of training data diversity and synthetic data on SLM performance. The default Slam recipe does not use diverse data (only Libri-light and LibriSpeech), but uses the synthetic sTinyStories data.
Diverse Data. We begin by examining how dataset diversity impacts model performance. Many leading speech datasets, such as those based on audiobooks (Panayotov et al., 2015; Kahn et al., 2020), consist of relatively clean, single-speaker recordings within a specific content domain. To introduce greater diversity in speaking styles and content, we curate additional datasets, including VoxPopuli Wang et al. (2021b), Tedlium Hernandez et al. (2018), PeopleSpeech Galvez et al. (2021), and SWC Baumann et al. (2018). For all mentioned datasets, we use the official data cleaning and preprocessing scripts when available. Specifically, for Libri-light, we apply the official Voice Activity Detection model to remove silences and generate smaller audio segments. To evaluate the impact of dataset diversity, we compare the performance of SLMs trained using our best training recipes using a subset of LibriSpeech and Libri-light against all curated datasets. This comparison is conducted for both OPT-$125$M, which processes a large number of tokens during training, and Qwen-$0.5$B, which encounters significantly less data due to model size. Results are summarised in Table 1. We observe that dataset diversity has an overall negative effect on model performance. We hypothesise this is due to the models struggling in modelling rich and complex audio under such low compute resources.
Figure 5: Analysing the optimal part of the 24 hour compute budget that should be used for DPO, with the rest used for pre-training.
Compute (GPU days) Parameters sBLIMP $\uparrow$ sStoryCloze $\uparrow$ tStoryCloze $\uparrow$ GenPPL $\downarrow$ Auto-BLEU $\downarrow$
TWIST-350M Hassid et al. (2024) 40*V100 305M 56.20 - - 137.3 3.46
TWIST-1.3B Hassid et al. (2024) 160*V100 1B 57.00 52.4 70.6 131.8 3.20
TWIST-7B Hassid et al. (2024) ? 7B 59.00 55.3 74.1 93.7 3.06
TWIST-13B Hassid et al. (2024) ? 13B 59.20 55.4 76.4 - -
Scaled Optimal Cuervo and Marxer (2024) ? 823M 61.3 56.7 78.0 - -
Predicted Optimal Cuervo and Marxer (2024) 1*A5000 78M 56.85 54.09 70.49 - -
TWIST-350M (Original recipe) 1*A5000 305M 51.52 $\pm$ .19 53.65 $\pm$ .57 68.80 $\pm$ .47 259.2 $\pm$ 6.7 3.26 $\pm$ .46
TWIST-350M + sTinyStories 1*A5000 305M 51.21 $\pm$ .26 54.17 $\pm$ .54 72.40 $\pm$ .18 159.0 $\pm$ 6.0 4.18 $\pm$ .24
Slam (-DPO) (ours) 1*A5000 358M 56.45 $\pm$ .17 55.59 $\pm$ .30 78.01 $\pm$ .27 88.3 $\pm$ 1.0 3.47 $\pm$ .17
Slam (ours) 1*A5000 358M 58.86 $\pm$ .20 58.04 $\pm$ .51 82.04 $\pm$ .21 62.8 $\pm$ 4.1 3.88 $\pm$ .11
Table 2: Comparing slamming to leading SLMs, and predicted optimal performance for the compute. We also consider TWIST-$350$M using our code and compute budget, but with the original training recipe. $\pm$ indicates distance to min/max of $3$ seeds.
Synthetic Data. Recent studies have highlighted the potential of synthetic data generated through Text-to-Speech (TTS) (Cuervo and Marxer, 2024) or direct text-to-unit conversion (Zeng et al., 2024). Hence, we examine the impact of including synthetically generated speech within our constrained compute setup. To do so, we synthesised the TinyStories dataset (Eldan and Li, 2023) using a single-speaker TTS model (Wang et al., 2021a), as it is computationally efficient. Additionally, prior research has shown that HuBERT units largely remove speaker information (Maimon and Adi, 2023). TinyStories has been demonstrated to enhance text LM performance and improve SLMs (Cuervo and Marxer, 2024). Results are presented in Table 1. Results indicate that incorporating such synthetic data into the training data-mix significantly boosts both modelling and generative performance metrics, across all evaluated setups. We also consider adding the synthetic data to the original TWIST recipe, and the results in the bottom of Table 2 suggests that while this helps with semantic metrics, it is far from enough without other optimisations we introduced.
We see that across all datasets, and specifically with our best mixture Libri-Light, LibriSpeech and sTinyStories, Qwen-$0.5$B outperforms OPT-$125$M so we continue with it to the final stages.
4.3 Text Interleaving
Several recent SLMs combine both speech and text modalities, either predicting both simultaneously (Défossez et al., 2024; Fang et al., 2024; Xie and Wu, 2024) or training on interleaved data (Nguyen et al., 2025; Zeng et al., 2024). Beyond enhancing cross-modal abilities, this has been shown to improve the semantic capabilities of SLMs, even in speech-only evaluations. Building on these studies, we investigate whether speech-text interleaving can enhance semantic ability in speech-only tasks, even under strict computational constraints.
For this we use Whisper-large-v3-turbo to get aligned transcriptions of our data, except sTinyStories for which we get alignment from the TTS. We follow Zeng et al. (2024) by selecting speech spans with length from a Poisson distribution with $\lambda=10$ totalling $30\%$ of the interleaved data. Following Nguyen et al. (2025) we train with balanced batches regarding token count between text data, speech data and interleaved data. We use a subset of RedPajama (Weber et al., 2024) filtered by Gopher (Rae et al., 2021) rules as our text data.
We find that under our setup performance is notably worse than speech-only Slam in all metrics. We hypothesise that in the slamming setup the larger vocabulary (leading to more parameters), and fewer speech tokens led to under-trained models. We leave for future work to find the minimal compute budget to benefit from text-interleaving or to consider other efficient approaches.
GPUs Params Num tokens sBLIMP $\uparrow$ sStoryCloze $\uparrow$ tStoryCloze $\uparrow$ GenPPL $\downarrow$ Auto-BLEU $\downarrow$
Speech only pre-training
GSLM Lakhotia et al. (2021) 8*V100 100M 1B 54.2 53.3 66.6 $\emptyset$ $\emptyset$
SyllableLM Baade et al. (2024) 4*A40 300M 16B 63.7 $\emptyset$ 75.4 $\emptyset$ $\emptyset$
TWIST-350M Hassid et al. (2024) 8*V100 305M 10.8B 56.20 $\emptyset$ $\emptyset$ 137.3 3.46
TWIST-1.3B Hassid et al. (2024) 32*V100 1B 10.8B 57.00 52.4 70.6 131.8 3.20
TWIST-7B Hassid et al. (2024) 32*V100 7B 36B 59.00 55.3 74.1 93.74 3.06
TWIST-13B Hassid et al. (2024) 32*V100 13B 36B 59.20 55.4 76.4 $\emptyset$ $\emptyset$
Scaled Optimal Cuervo and Marxer (2024) $\emptyset$ 823M 82B 61.3 56.7 78.0 $\emptyset$ $\emptyset$
Moshi Défossez et al. (2024) ?*H100 7B ? 58.9 58.7 81.8 $\emptyset$ $\emptyset$
SpiritLM Nguyen et al. (2025) 64*A100 7B 100B 58.0 54.8 72.9 $\emptyset$ $\emptyset$
Joint speech-text pre-training / preference optimisation
Zeng et al. (2024) $\emptyset$ 9B $\sim$1T $\emptyset$ 62.4 82.9 $\emptyset$ $\emptyset$
Moshi Défossez et al. (2024) ?*H100 7B $\sim$720B 58.8 60.8 83.0 $\emptyset$ $\emptyset$
SpiritLM Nguyen et al. (2025) 64*A100 7B 100B 58.3 61.0 82.9 $\emptyset$ $\emptyset$
AlignSLM-1.3B Lin et al. (2024) 64*A100 1B 10.8B + $\sim$158B 59.8 55.0 80.0 $\emptyset$ $\emptyset$
AlignSLM-7B Lin et al. (2024) 64*A100 7B 36B + $\sim$158B 62.3 61.1 86.8 $\emptyset$ $\emptyset$
Slam (-DPO) 2*A100 358M 16.7B 58.53 58.15 80.71 67.3 3.25
Slam 1*A5000 358M 1.4B + 5M 58.86 58.04 82.04 62.8 3.88
Slam (scaled) 2*A100 358M 16.7B + 9M 61.11 61.30 84.18 46.6 3.75
Table 3: Analysing the effect of scaling up compute for Slam. Number tokens refers to total, not necessarily unique, tokens used for training (estimated from the provided information). We separately mark DPO tokens with a +.
4.4 Synthetic Data Preference Optimisation
Preference optimisation methods have been shown to enhance the performance of text LLMs Ouyang et al. (2022) and, more recently, SLMs Lin et al. (2024). With preference optimisation, we aim to train our model to generate outputs that better align with a specified reward function or preference set.
We evaluate how preference optimisation affects SLM performance while considering our constrained computational budget. Using an off-policy approach with pre-generated preference data, we apply DPO to enhance training efficiency. Specifically, we synthetically generate the SWAG (Zellers et al., 2018) text corpus for evaluating semantic knowledge. SWAG consists of text prefixes paired with multiple possible suffixes, where only one is semantically plausible. For preference data, we use the first sentence as the prompt, the correct suffix as the positive continuation, and a randomly chosen incorrect suffix as the rejected continuation. To ensure quality, we filter out samples with repetitive patterns, identified by an auto-BLEU score above $0.3$. We generate all recordings using Kokoro TTS (Hexgrad, 2025), incorporating four speakers (two male and two female), evenly split between British and American accents. This process results in a total of $47$k SWAG preference pairs.
For DPO we use $\beta=0.1$ (see Appendix A for full hyperparameters). In initial tests, we observe that after DPO training, the model shows increased likelihood at the cost of repeated patterns, a known issue with DPO (Lanchantin et al., 2025). To address this, we apply a repetition penalty with a factor of $1.1$, following the approach of Keskar et al. (2019), and find that it helps mitigate the problem. Future work could explore alternative solutions, such as proposed by Lanchantin et al. (2025).
We begin by examining how the allocation of budget for DPO impacts performance, particularly when it comes at the cost of a shorter pre-training phase. Figure 5 depicts the results. We observe significant improvements across all metrics when applying DPO for at least $30$ minutes compared to not using DPO at all. However, allocating a higher proportion of the budget to DPO does not yield further gains and can even degrade model performance. Thus we stick to $30$ minutes out of $24$ hours for DPO, using the rest for pre-training.
5 Final Recipe
Building on these empirical findings, we develop the final Slam recipe. Using it, we train SLMs based on Qwen2.52.52.52.5-$0.5$B. We then compare Slam to the TWIST model family across various sizes: $350$M, $1.3$B, $7$B, and $13$B. We also present results for TWIST-$350$M using our computational constraints but following TWIST’s original training recipe, along with our synthetic data. Finally, we report results for the top-performing model from  Cuervo and Marxer (2024), including their predicted optimal performance under our compute budget based on SLM scaling laws. Results are reported in Table 2. The results indicate that Slam delivers performance that is either superior or on par with baseline models while requiring significantly fewer computational resources (e.g., a single A5000500050005000 for a day compared to $160$ days on a V100100100100).
6 Increasing Compute
Similarly to Geiping and Goldstein (2023), we analyse whether the proposed approach holds well also in increased compute budget. We opt for $48$ hours on $2$ A100100100100 GPUs as a reasonable academic budget for larger scale tests, and represents $\sim 10$ times more compute than the Slamming setting. We use exactly the same Slam recipe for more steps, and increase the batch size times $2$. We provide the full results in Table 3. We note that the performance continues to improve across all metrics, also outperforming methods which have far larger compute scales. We note that DPO training on synthetic data for $2$ epochs, notably boosts performance.
7 Conclusion
In this work we show that training high quality SLMs with a very modest compute budget, is feasible. We give these main guidelines: (i) Do not skimp on the model - not all model families are born equal and the TWIST initialisation exaggerates this, thus it is worth selecting a stronger / bigger text-LM even if it means less tokens. we found Qwen2.52.52.52.5 to be a good choice; (ii) Utilise synthetic training data - pre-training on data generated with TTS helps a lot; (iii) Go beyond next token prediction - we found that DPO boosts performance notably even when using synthetic data, and as little as $30$ minutes training massively improves results; (iv) Optimise hyper-parameters - as researchers we often dis-regard this stage, yet we found that tuning learning rate schedulers and optimising code efficiency can improve results notably. We hope that these insights, and open source resources will be of use to the research community in furthering research into remaining open questions in SLMs.
Limitations
While the SLMs trained under Slamming compute budget performed notably well compared to other SLMs trained with much more compute they might perform less well in other areas. For instance, evaluating their abilities on acoustic or prosodic elements as in SALMon Maimon et al. (2024) could show further challenges of low resource settings.
Furthermore, we focus in this study on the well used HuBERT Hsu et al. (2021) model as a tokeniser, and while we do not make any adjustments specifically for it, future work might wish to investigate our cramming approach with new tokenisers, such as Mimi Défossez et al. (2024) and SylBoost Baade et al. (2024).
Ethical Statement
The broader impact of this study is, as in any generative model, the development of a high quality and natural speech synthesis. We hope that allowing training SLMs under low-resource settings, and open sourcing resources to aid this goal, will have a positive impact on inclusivity and accessibility of SLM research beyond well funded labs.
Appendix A Full Slam Recipe
We provide below the full training recipe, including hyperparameters for the best, Slam recipe. In Table  4 we see the Slam (-DPO) pre-training recipe and in Table  5 we see the Slam DPO training recipe.
Table 4: Slam (-DPO) Pre Training Recipe
Parameter Value
Text Base Model Qwen2.5-0.5B
TWIST initialisation True
Data Librilight + Librispeech + sTinyStories
Train Time $23.5$ hours $\simeq 17625$ steps
RoPE theta $10000$
Context length $1024$
Per device Batch Size $8$
Gradient Accumulation $16$
Base Learning Rate $31e-$ 3
Warmup Ratio $1\%$
Optimizer AdamW
Learning Rate Scheduler cosine with min $55e-$ 5
Max Grad Norm $0.5$
Dtype bfloat16
Table 5: Slam DPO Training Recipe
Parameter Value
Initial Model Slam (-DPO)
Data SpokenSwag with auto-bleu smaller than 0.3
Train Time $0.5$ hour $\simeq 813$ steps
RoPE theta $10000$
Context length $1024$
Per device Batch Size $4$
Gradient Accumulation $16$
Base Learning Rate $55e-$ 5
Optimizer AdamW
Learning Rate Scheduler inverse sqrt
Max Grad Norm $0.5$
Dtype bfloat16
DPO $\beta$ $0.1$
Appendix B Model Sizes
Table 6: Model names and parameter counts after changing vocabulary to speech only units (500).
Model Name Number of Params
MobileLLM-125M Liu et al. (2024b) 106,492,608
MobileLLM-350M Liu et al. (2024b) 315,117,120
OPT-125M Zhang et al. (2022) 87,015,936
OPT-350M Zhang et al. (2022) 305,714,176
QWEN2.5-0.5B Yang et al. (2024a) 358,347,904
SmolLM2-135M Allal et al. (2025) 106,492,608
SmolLM2-360M Allal et al. (2025) 315,117,120
Pythia-160M Biderman et al. (2023) 85,827,072
Pythia-410M Biderman et al. (2023) 303,339,520
As mentioned, we use the original names of the text LMs used for clarity and consistency, but note that the actual parameter counts after resizing the vocabulary to speech-units only can be very different. In Table 6 we provide an extensive list of models and sizes.
Appendix C Dataset Statistics
We use and synthesise several datasets. In this section we give exact details of number of samples, splits used, domains etc.
For pre-training we use Libri-Light Kahn et al. (2020) and LibriSpeech Panayotov et al. (2015). For Libri-Light we randonly select one percent of samples as validation, whereas for LibriSpeech we use the original dev-clean and dev-other splits. Both of these datasets are English speech only, focused in the audio-book domain. We also synthesise sTinyStories for pre-training which consists of synthetically generated English short stories. We use the official train split for training. Full dataset sizes are in Table 7.
We also investigate diverse datasets for pre-training: SWC Baumann et al. (2018), Tedlium Hernandez et al. (2018), PeopleSpeech Galvez et al. (2021) and VoxPopuli Wang et al. (2021b). We only take English subsets for all datasets, yet they can still contain diverse accents. These datasets are in the following domains SWC - read Wikipedia articles, Tedlium - short lectures, PeopleSpeech - diverse data including many local council gatherings etc, VoxPopuli - from European Parliament meetings. For SWC specifically, we use the text alignment to create chunks, remove silence from the audio and remove mis-aligned chunks. We use full training splits where provided, otherwise splitting 99% for training. The dataset sizes are described in 7.
Table 7: Dataset train set sizes that we use.
Dataset Number of Hours Number of Tokens
Libri-Light Kahn et al. (2020) $50K$ $3.5B$
LibriSpeech Panayotov et al. (2015) $960$ $67M$
SWC Baumann et al. (2018) $750$ $19M$
Tedlium Hernandez et al. (2018) $1.6K$ $110M$
PeopleSpeech Galvez et al. (2021) $7K$ $480M$
VoxPopuli Wang et al. (2021b) $24K$ $1.64B$
sTinyStories $30K$ $2.2B$
For DPO we synthesise SpokenSwag based on the SWAG Zellers et al. (2018) dataset. We use only the official train set and filter only the gold standard labels. We end up with 47k sample pairs which end up to be $\sim 4.5M$ tokens.
Appendix D AI Tool Usage
AI based tools may have been used in writing parts of the code for this study, or para-phrasing some of the writing within the paper, yet all the content was thoroughly checked by the authors, with these only being used as assistive tools.